OpenStudy (anonymous):

Please Help, Write an equation parallel to the line and shifted 4 units up from this line 2x-y =8 Please show me step by step so I can do it on my own

OpenStudy (misty1212):

\[y=2x-8\]shift up four and you will have \[y=2x-4\] or \[2x-y=4\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

ok confused

OpenStudy (triciaal):

parallel lines have the same slope. the intercept will be different shift up means add units to the y value

OpenStudy (triciaal):

y = mx + b given line m = 2 and b = -8 parallel line m = 2 and b = shift 4 up = -8 + 4 = -4 so the parallel line y= 2x - 4

OpenStudy (triciaal):

looks like this